# LogicGame: Benchmarking Rule-Based Reasoning Abilities of Large Language Models

### v3

@@ 1 Introduction
In this paper, we introduce LogicGame, a novel benchmark crafted to evaluate the comprehensive rule understanding, planning, and execution capabilities of LLMs. LogicGame offers a set of carefully designed rule-based reasoning games. Each game contains a series of rules the model must follow to find the solution involving single or multiple steps. During the data construction process, we ensure that all the problems remain unavailable on the Internet to prevent data leakage. LogicGame covers two main scenarios: execution and planning, each divided into several sub-categories. Execution problems include tasks related to string data manipulation, where models handle string data transformations, as well as arithmetic operations and manipulations, focusing on mathematical computations and sequential execution. Planning games encompass math puzzles, which require solving complex mathematical problems, and pure logic puzzles, involving abstract reasoning without numerical computation. Through these varied games, LogicGame aims to comprehensively evaluate the rule-based reasoning capabilities in LLMs.
@@ 4.1 Experimental Setup
We evaluate 11 popular LLMs. Closed-source models included versions from Claude [1] and GPT-4 [26] series. Open-source models encompassed LLaMA 3 [33], Qwen [2], GLM [12], Mistral [17], and InternLM [32] variants.
In the inference stage, we set temperature to 0, ensuring deterministic outputs. The maximum token number is set to 2048. Other parameters are set as their default values.
4.2 Main Results and Analysis
The average performance of 11 LLMs on LogicGame across our Chinese and English bilingual dataset, measured by AP-Acc, is presented in Table 2. The results indicate that reasoning remains a significant challenge for LLMs. Even the strongest model can’t reach an overall score of 21%. As the difficulty level increases, the performance of LLMs drops significantly, with many models dropping to 0% score at Level 3, especially in Planning tasks. This suggests that even advanced LLMs struggle with complex reasoning tasks.
The performance drop from Level 0 to Level 3 is not uniform across models. Some models (e.g. claude-3.5-sonnet) show a more gradual decline, while others drop off sharply after Level 1. This may suggest that some models have better consistency across task complexities.
Interestingly, the performance of models in Execution and Planning tasks varies. Some top-performing models, such as claude-3.5-sonnet, demonstrates superior performance in Execution compared to Planning (22.50% vs. 16.67%). Conversely, llama-3-70b-chat excels in Planning over Execution (10.19% vs. 5.56%). Even within model families, the relative performance differences are evident. In the GPT family, gpt-4-turbo-0409 outperforms gpt-4o in Planning tasks (16.21% vs. 13.43%), while gpt-4o shows better performance in Execution tasks (19.44% vs. 15.83%).
@@ 4.2 Main Results and Analysis
Model & Execution(%) Planning(%) Overall(%)
@@
claude-3.5-sonnet 55.56 17.78 11.11 5.56 22.50 44.44 16.67 1.85 3.70 16.67 20.93
gpt-4o 43.33 16.67 12.22 5.56 19.44 33.33 16.67 3.70 0.00 13.43 18.43
gpt-4-turbo-0409 38.89 14.44 6.67 3.33 15.83 46.30 14.81 3.70 0.00 16.21 17.04
glm-4-plus 34.44 11.11 7.78 1.11 13.61 35.19 16.67 1.85 0.00 13.43 14.64
qwen2-72b-instruct 31.11 10.00 3.33 0.00 11.11 33.33 11.11 0.00 0.00 10.93 12.01
llama-3-70b-chat 16.67 5.56 0.00 0.00 5.56 31.48 9.26 0.00 0.00 10.19 8.28
claude-3-haiku 14.44 6.67 4.44 0.00 6.39 20.37 3.70 0.00 0.00 6.02 6.80
glm-4-9b 18.89 1.11 0.00 0.00 4.72 18.52 3.70 0.00 0.00 5.56 5.76
internlm-2.5-7b-chat 6.67 2.22 1.11 0.00 2.50 18.52 0.00 1.85 0.00 5.10 4.28
qwen2-7b-instruct 10.00 2.22 0.00 0.00 3.06 14.81 1.85 0.00 0.00 4.17 3.29
mistral-7b-instruct 3.33 0.00 0.00 0.00 0.83 5.56 0.00 0.00 0.00 1.39 1.97
llama-3-8b-chat 1.11 0.00 0.00 0.00 0.28 5.56 1.85 0.00 0.00 1.85 0.81
@@ 4.3 Few-shot Results
In the "Execution" category, models demonstrate notable improvements in accuracy with increased shot contexts demonstrated in Figure 3. Specifically, stronger models (as indicated in Table 2) like gpt-4o, qwen2-72b-instruct shows a greater increase in the AP-Acc score when transitioning from 0-shot to 1-shot and 2-shot settings than weaker ones, indicating enhanced execution accuracy with more contextual information. However, the effects of 1-shot and 2-shot settings vary across models Performance variations by difficulty levels, as shown in Figure 4, indicate that models benefit most from 1-shot and 2-shot contexts at Level 0. And in general, the influence of shot contexts diminishes as the difficulty level increases. This consistency suggests that simpler tasks (Level 0) allow models to leverage additional context more effectively, enhancing their execution capabilities across the board.
@@ Appendix C Additional evaluation metrics and error analysis
claude-3.5-sonnet 0.00 0.00 0.00 0.00 22.51 15.80 19.15
gpt-4o 1.67 0.45 0.00 0.00 20.07 13.28 16.67
gpt-4-turbo-0409 0.84 0.00 0.56 0.00 16.12 15.32 15.72
glm-4-plus 2.50 0.93 0.28 0.00 14.00 13.55 14.93
qwen2-72b-instruct 0.28 0.45 0.00 0.00 11.16 10.98 11.07
llama-3-70b-chat †16.95 †15.16 0.28 †0.88 7.89 13.10 10.50
claude-3-haiku 0.00 0.00 0.00 0.00 6.41 5.70 6.06
glm-4-9b 9.72 1.84 0.00 0.00 5.49 5.56 5.23
llama-3-8b-chat ‡62.78 ‡55.96 0.00 0.45 2.19 6.97 4.58
internlm-2.5-7b-chat 12.50 0.90 0.00 0.00 3.07 5.15 4.11
qwen2-7b-instruct 0.84 0.47 †3.61 0.00 3.35 4.14 3.74
mistral-7b-instruct 16.67 2.71 ‡5.28 †2.73 1.33 1.53 1.43
The metrics presented in Table 4 provide a comprehensive assessment of the model’s error frequency. The NIJ-Acc metric evaluates the model’s overall accuracy. In this study, the llama-3-8b-chat model shows poor instruction-following capabilities, while the mistral-7b-instruct model performs poorly across all error metrics. Conversely, stronger models tend to perform more stably as observed in these metrics. Overall error probabilities (IFError, JSError) are low across most models, resulting in minimal variations in NIJ-Acc scores compared to the average scores in Table 2, with only minor rank adjustments among similarly performing models.
@@ Appendix E Problems all models fail
Figure 9 categorizes the five areas where all models exhibit the poorest performance, presenting the average AP-ACC scores for each category in a heat map. The horizontal axis corresponds to the model capabilities as outlined in Table 2. This figure highlights that models generally struggle most in two sub-categories within the execution scenario, particularly with ’Reversi’, where many models score close to zero. Conversely, in planning scenarios such as ’Constrained Linear Arrangement’, there is a slight variation in performance across different models.

### v4

@@ 1 Introduction
In this paper, we introduce LogicGame, a novel benchmark crafted to evaluate the comprehensive rule understanding, planning, and execution capabilities of LLMs. LogicGame offers a set of carefully designed rule-based reasoning games. Each game contains a series of rules the model must follow to find the solution involving single or multiple steps. During the data construction process, we ensure that all the problems remain unavailable on the Internet to prevent data leakage. LogicGame covers two main scenarios: execution and planning, each divided into several sub-categories. Execution problems include tasks related to string data manipulation, where models handle string data transformations, as well as arithmetic operations and manipulations, focusing on mathematical computations and sequential execution. Planning games encompass math puzzles, which require solving complex mathematical problems, and pure logic puzzles, involving abstract reasoning without numerical computation. Through these varied games, LogicGame aims to comprehensively evaluate the rule-based reasoning capabilities in LLMs.111We have released the dev set and the whole set of the LogicGame, and created a leaderboard on https://github.com/Hypatiaalegra/LogicGame-Data. You can refer to https://www.codabench.org/competitions/4140/ for a fair and fast evaluation of the LogicGame.
@@ 4.1 Experimental Setup
We evaluate 14 popular LLMs. Closed-source models included versions from Claude [1] and GPT [26] series. Open-source models encompassed LLaMA 3 [33], Qwen [2], GLM [12], Mistral [17], and InternLM [32] variants. In the inference stage, we set temperature to 0, ensuring deterministic outputs. The maximum token number is set to 2048. Other parameters are set as their default values.
@@
Model & Execution Planning Overall
@@
o1-preview 82.22 66.67 48.89 42.22 60.00 64.52 58.06 35.48 32.26 47.58 54.93
o1-mini 77.78 57.78 48.89 42.22 56.67 74.19 58.06 35.48 12.90 45.16 51.97
claude-3-5-sonnet 68.89 40.00 22.22 15.56 36.67 48.39 25.81 6.45 3.23 20.97 30.26
gpt-4o 62.22 31.11 13.33 13.33 30.00 51.61 22.58 9.68 6.45 22.58 26.97
gpt-4-turbo-0409 60.00 15.56 15.56 11.11 25.56 67.74 22.58 9.68 3.23 25.81 25.66
glm-4-plus 42.22 26.67 17.78 6.67 23.33 48.39 19.35 6.45 3.23 19.36 21.71
qwen2-72b 53.33 20.00 15.56 2.22 22.78 45.16 16.13 3.23 3.23 16.94 20.39
llama-3-70b 42.22 11.11 6.67 2.22 15.56 25.81 6.45 0.00 0.00 8.07 12.50
claude-3-haiku 31.11 4.44 2.22 0.00 9.44 32.26 6.45 0.00 0.00 9.68 9.54
glm-4-9b 24.44 8.89 2.22 0.00 8.89 19.35 3.23 0.00 0.00 5.65 7.57
internlm2-5-7b 13.33 4.44 0.00 0.00 4.44 16.13 3.23 0.00 0.00 4.84 4.61
llama-3-8b 11.11 2.22 0.00 0.00 3.33 9.68 6.45 0.00 0.00 4.03 3.62
mistral-7b 4.44 0.00 0.00 0.00 1.11 19.35 3.23 0.00 0.00 5.65 2.96
qwen2-7b 4.44 0.00 0.00 0.00 1.11 16.13 3.23 0.00 0.00 4.84 2.63
hline1, Z = 1pt, solid,
hline2, 3,
cell11 = r = 2m,
cell12 = c = 5c,
cell17 = c = 5c,
cell112 = r = 2m,
vline2, 6, 7, 11, 12,
Model & Execution Planning Overall
Level 0 Level 1 Level 2 Level 3 Avg. Level 0 Level 1 Level 2 Level 3 Avg.
o1-preview 71.11 57.78 48.89 51.11 57.22 70.97 54.84 41.94 22.58 47.58 53.29
o1-mini 73.33 46.67 44.44 46.67 52.78 70.97 48.39 38.71 22.58 45.16 49.67
claude-3-5-sonnet 46.67 40.00 33.33 13.33 33.33 64.52 16.13 6.45 6.45 23.39 29.28
gpt-4o 57.78 37.78 31.11 13.33 35.00 48.39 19.35 3.23 3.23 18.55 28.29
gpt-4-turbo-0409 42.22 20.00 17.78 8.89 22.22 51.61 12.90 9.68 3.23 19.36 21.05
glm-4-plus 31.11 15.56 17.78 6.67 17.78 48.39 9.68 9.68 3.23 17.75 17.76
qwen2-72b 28.89 4.44 0.00 0.00 8.33 22.58 6.45 6.45 3.23 9.68 8.88
glm-4-9b 22.22 6.67 2.22 2.22 8.33 22.58 6.45 0.00 0.00 7.26 7.89
internlm2-5-7b 22.22 4.44 4.44 0.00 7.78 12.90 3.23 0.00 0.00 4.03 6.25
claude-3-haiku 8.89 8.89 0.00 0.00 4.45 22.58 0.00 0.00 0.00 5.65 4.93
llama-3-70b 8.89 8.89 8.89 0.00 6.67 3.23 3.23 0.00 0.00 1.62 4.61
mistral-7b 22.22 0.00 0.00 0.00 5.56 9.68 0.00 0.00 0.00 2.42 4.28
qwen2-7b 4.44 2.22 0.00 0.00 1.67 6.45 0.00 0.00 0.00 1.61 1.64
llama-3-8b 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00 0.00
4.2 Main Results and Analysis
Table 3 presents the performance of 14 LLMs on LogicGame across our Chinese and English bilingual dataset, measured by AP-Acc. The o1-preview model leads with 54.93% and 53.29% overall accuracy for Chinese and English respectively, closely followed by o1-mini. These results underscore the persistent challenge of reasoning for LLMs, as even top performers barely exceed 50% accuracy. The substantial performance gap, ranging from over 50% to below 5%, not only highlights the varying capabilities of current LLMs in complex reasoning tasks, but also emphasizes that despite recent advancements, logical reasoning remains a significant hurdle for most language models.
The performance drop from Level 0 to Level 3 is not uniform across models. Some models (e.g. o1-mini) show a more gradual decline, while others drop off sharply after Level 1. This may suggest that some models have better consistency across task complexities.
Interestingly, the performance of models in Execution and Planning tasks varies. Some top-performing models, such as o1-preview, demonstrates superior performance in Execution compared to Planning. Conversely, llama-3-70b-chat excels in Planning over Execution. Even within model families, the relative performance differences are evident. In the GPT family, gpt-4-turbo-0409 outperforms gpt-4o in Planning tasks, while gpt-4o shows better performance in Execution tasks.
@@ 4.3 Few-shot Results
In the "Execution" category, models demonstrate notable improvements in accuracy with increased shot contexts demonstrated in Figure 4. Specifically, stronger models (as indicated in Table 3(a) and Table 3(b)) like gpt-4o, qwen2-72b-instruct shows a greater increase in the AP-Acc score when transitioning from 0-shot to 1-shot and 2-shot settings than weaker ones, indicating enhanced execution accuracy with more contextual information. However, the effects of 1-shot and 2-shot settings vary across models Performance variations by difficulty levels, as shown in Figure 5, indicate that models benefit most from 1-shot and 2-shot contexts at Level 0. And in general, the influence of shot contexts diminishes as the difficulty level increases. This consistency suggests that simpler tasks (Level 0) allow models to leverage additional context more effectively, enhancing their execution capabilities across the board.
@@ Appendix C Additional evaluation metrics and error analysis
o1-preview 0.56 0.00 0.00 0.00 57.54 47.58 53.47
o1-mini 0.56 0.81 0.56 0.00 53.37 45.53 50.17
claude-3.5-sonnet 0.56 0.00 0.00 0.00 33.52 23.39 29.37
gpt-4o 1.11 0.00 0.00 1.61 35.39 18.85 28.67
gpt-4-turbo-0409 1.11 0.00 0.00 †2.42 22.47 19.83 21.40
glm-4-plus 5.00 9.68 1.11 0.00 19.64 18.93 19.22
qwen2-72b-instruct 30.00 16.94 0.56 0.81 12.00 11.76 11.89
llama-3-70b-chat †52.22 ‡60.48 0.56 0.00 14.12 4.08 10.45
glm-4-9b 21.67 10.48 2.22 0.81 10.95 8.18 9.72
internlm-2.5-7b-chat 15.56 9.68 2.22 0.00 9.46 4.46 7.31
mistral-7b-instruct 19.44 14.52 ‡13.33 †2.42 8.26 2.91 5.80
claude-3-haiku 0.00 0.00 1.11 †2.42 4.49 5.49 5.02
qwen2-7b-instruct 2.22 0.81 †11.67 ‡3.23 1.94 1.68 1.82
llama-3-8b-chat ‡75.00 †58.87 0.00 0.81 0.00 0.00 0.00
Execution Planning Execution Planning Execution Planning Avg.
o1-preview 0.00 0.00 0.00 0.00 60.00 47.58 54.93
o1-mini 0.56 0.81 0.56 0.00 56.67 45.16 51.97
claude-3.5-sonnet 0.00 0.00 0.00 0.00 36.67 20.97 30.26
gpt-4o 0.56 0.00 0.00 0.00 30.17 22.58 27.06
gpt-4-turbo-0409 0.56 1.61 0.00 1.61 25.70 26.67 26.09
glm-4-plus 10.00 8.06 0.56 0.00 26.09 21.05 24.00
qwen2-72b-instruct 3.33 1.61 0.56 1.61 23.70 17.50 21.16
llama-3-70b-chat 12.78 †13.71 0.00 0.81 17.83 9.43 14.45
claude-3-haiku 3.89 0.81 0.00 0.81 9.83 9.84 9.83
glm-4-9b 25.00 †13.71 1.11 0.81 12.03 6.60 9.62
llama-3-8b-chat ‡40.56 ‡28.23 0.00 0.00 5.61 5.62 5.61
internlm-2.5-7b-chat 17.78 1.61 1.67 ‡5.65 5.52 5.22 5.38
mistral-7b-instruct †36.67 11.29 ‡7.22 †4.84 1.98 6.73 4.39
qwen2-7b-instruct 3.89 2.42 †2.22 4.03 1.18 5.17 2.81
The metrics presented in Table 3 and Table 4 provide a comprehensive assessment of the model’s error frequency in en and zh version respectively. The NIJ-Acc metric evaluates the model’s overall accuracy. In this study, the llama-3-8b-chat and llama-3-70b-chat model shows consistent poor instruction-following capabilities in both versions, while the mistral-7b-instruct model performs poorly in the metric of json format. Conversely, stronger models tend to perform more stably as observed in these metrics. Overall error probabilities (IFError, JSError) are low across most models, resulting in minimal variations in NIJ-Acc scores compared to the average scores in Table LABEL:tab:main_result, with only minor rank adjustments among similarly performing models.
@@ Appendix E Problems all models fail
Figure 10 categorizes the five areas where all models exhibit the poorest performance, presenting the average AP-ACC scores for each category in a heat map. The horizontal axis corresponds to the model capabilities as outlined in Table 3(b) and Table 3(a). This figure highlights that models generally struggle most in two sub-categories within the execution scenario, particularly with ’Reversi’, where many models score close to zero except for o1-mini and o1-preview models. Conversely, in planning scenarios such as ’Constrained Linear Arrangement’, there is a slight variation in performance across different models.